# Non-toric brane webs, Calabi–Yau 3-folds, and 5d SCFTs

### v1

@@ 4.4. Flops and pushing 7-branes
obtained from $\pi:(\mathcal{Y},\mathcal{D},\mathcal{L})\rightarrow\mathbb{C}$ by a sequence of M1 flops. Indeed, pushing a 7-brane $x_{e,i}$ across $v$ in $W_{v}$ as in Case II-(4.15) with $w_{-e}\geq 1$ amounts to flopping the $(-1)$-curve $E_{e,i}^{\prime}$ as in §4.1. On the other hand, pushing a 7-brane $x_{e,i}$ across $v$ in $W_{v}$ as in Case II-(4.15) with $w_{-e}=\mu=0$, or as in Case III-(4.16) does not change the degeneration, but only the toric model of the log Calabi–Yau surface $(Y_{0}^{v},\partial Y_{0}^{v},L_{0}^{v})$, as in the geometric description of Hanany–Witten moves in §4.1.
@@
Since $W$ is consistent, it follows from Lemma 4.13 that there exists a degeneration of log Calabi–Yau surfaces $\pi^{\prime}:(\mathcal{Y}^{\prime},\mathcal{D}^{\prime},\mathcal{L}^{\prime})\rightarrow\mathbb{C}$, obtained from $\pi:(\mathcal{Y},\mathcal{D},\mathcal{L})\rightarrow\mathbb{C}$ by a sequence of M1 flops, such that all line bundles $(L^{\prime})_{0}^{v}$ are nef. This sequence of M1 flops can be realized by a sequence of pushing operations of 7-branes to produce a modified web $W^{\prime}$. Indeed, as in the proof of Lemma 4.4, by [55, Proposition 3.27] (see also [21, Theorem 1]), any two toric models of a log Calabi–Yau surface are connected by a sequence of corner blow-ups/blow-downs and elementary cluster transformations, and it follows that any interior $(-1)$-curve can be realized after a finite sequence of pushing operations of 7-branes as the $(-1)$-curve $E_{e,i}^{\prime}$ corresponding to a pushed 7-brane $x_{e,i}^{\prime}$. ∎
@@ 4.5.1. Pushing a 7-brane through a 3-valent vertex
Let $W$ be the web of 5-branes with 7-branes obtained from $\overline{W}$ by requiring that all the $w_{3}$ 5-branes of direction $v_{3}$ end on the same 7-brane, see Figure 4.4.
[…]
Figure 4.4. Before and after pushing a 7-brane through a 3-valent vertex.
@@
It follows that we are in Cases I-II-III as in (4.14)-(4.15)-(4.16) if and only if $w_{3}^{2}>m$, $w_{3}^{2}=m$, and $w_{3}^{2}<m$ respectively. In particular, we are in Case II if and only if $w_{3}^{2}=m$, in which case the 7-brane can be pushed through the origin and is no longer attached to any 5-brane, as in Figure 4.4. The condition $w_{3}^{2}=m$ can be rephrased in terms of the lattice triangle $\overline{P}$ dual to $\overline{W}$. First note that the weights $w_{1},w_{2},w_{3}$ are the lattice lengths of the edges $e_{1},e_{2},e_{3}$ of $\overline{P}$ corresponding to the edges of $\overline{W}$. Denoting by $h$ the height of $\overline{P}$ with respect to $e_{3}$ the lattice distance between $e_{3}$ and the opposite vertex of $\overline{P}$, we have $m=w_{3}h$. It follows that the condition $w_{3}^{2}=m$ is equivalent to $w_{3}=h$.
We now describe the geometric interpretation of pushing the 7-brane through the origin $w_{3}=h$. Let $(\overline{Y},\overline{D},\overline{L})$ be the polarized toric surface corresponding to $\overline{W}$. The toric boundary divisor $\overline{D}$ has three irreducible toric divisors $\overline{D}_{1},\overline{D}_{2},\overline{D}_{3}$ corresponding to $e_{1},e_{2},e_{3}$, and the log Calabi–Yau surface $Y$ is obtained from $\overline{Y}$ by blowing up a smooth point $x_{3}$ of the toric boundary on $\overline{D}_{3}$, see Figure 4.5. Moreover, denoting by $E$ the exceptional divisor, we have $L=\overline{L}\otimes\mathcal{O}(-w_{3}E)$. Denote by $H$ the image in $Y$ of the strict transform $E^{\prime}$ of the fiber passing through $x_{3}$ of the $\mathbb{P}^{1}$-fibration on the toric blow-up of $\overline{Y}$ obtained by adding the ray $\mathbb{R}_{<0}v_{3}$ to the fan. One can check that $L\cdot H=h-w_{3}$. In particular, when $w_{3}=h$, we have $L\cdot H=0$ and so $L$ is nef. On the other hand, we have $H^{2}=0$ since $(E^{\prime})^{2}=-1$ and the toric divisor corresponding to $\mathbb{R}_{<0}v_{3}$ is contracted in $Y$. It follows that there is a fibration of $Y$ with fibers of class $H$. Finally, using that $L\cdot E=w_{3}>0$, one concludes that the map contracting all curves in $Y$ having zero-intersection with $L$ is a map $c:Y\rightarrow\mathbb{P}^{1}$. The fact that the surface $Y$ is contracted onto $\mathbb{P}^{1}$ is the geometric realization of the fact that, after pushing the 7-brane across the origin, the 3-valent vertex of the web disappears and the remaining 5-branes form a straight line (taking into account the monodromy cut), which can be viewed as the fan of $\mathbb{P}^{1}$.
[…]
Figure 4.5. The contraction map $c:Y\rightarrow\mathbb{P}^{1}$.
@@
[…]
Figure 4.6. Examples of lattice triangles with $w_{3}=h$.
@@
The left and middle lattice triangles in Figure 4.6 are elementary. The right triangle in Figure 4.6 satisfies $w_{3}=h$ but is not elementary.
@@ 4.5.2. Pushing a 7-brane through a 4-valent vertex
Let $\overline{W}$ be a web of 5-branes consisting of a single 4-valent vertex at the origin with edges of primitive directions of the form $\overline{v}_{1}$, $\overline{v}_{2}$, $\overline{v}_{3}$, $\overline{v}_{4}=-\overline{v}_{3}$. Let $w_{1},w_{2},w_{3},w_{4}$ be the weights of the edges, and denote $v_{i}=w_{i}\overline{v}_{i}$ for $1\leq i\leq 4$. By the balancing condition, we have $\sum_{i=1}^{4}v_{i}=0$. Let $W$ be the web of 5-branes with 7-branes obtained from $\overline{W}$ by requiring that all the 5-branes in direction $v_{3}$ all end on the same 7-brane, see Figure 4.7.
[…]
@@
Figure 4.7. Before and after pushing a 7-brane through a 4-valent vertex.
@@
Arguing as in §4.5.1, we obtain that, after moving the 7-brane through the origin, the number of 5-branes starting at the origin and ending on the 7-brane is $|\det(v_{2},\overline{v}_{3})|+w_{4}-w_{3}$. Note that, by the balancing condition, we have $|\det(v_{2},\overline{v}_{3})|=|\det(v_{1},\overline{v}_{3})|$. Hence, we are in Cases I-II-III as in (4.14)-(4.15)-(4.16) if and only if $w_{3}-w_{4}>|\det(v_{2},\overline{v}_{3})|$, $w_{3}-w_{4}=|\det(v_{2},\overline{v}_{3})|$, and $w_{3}-w_{4}<|\det(v_{2},\overline{v}_{3})|$ respectively. In particular, we are in Case II if and only if $w_{3}-w_{4}=|\det(v_{2},\overline{v}_{3})|$, in which case the 7-brane can be pushed through the origin and is no longer attached to 5-branes connected to the origin – see Figure 4.7. The condition $w_{3}-w_{4}=|\det(v_{2},\overline{v}_{3})|$ can be rephrased in terms of the lattice trapezoid $\overline{P}$ dual to $\overline{W}$. First note that the weights $w_{1},w_{2},w_{3},w_{4}$ are the weights of the edges $e_{1},e_{2},e_{3}$ of $\overline{P}$ corresponding to the edges of $\overline{W}$. Then, we have $|\det(v_{2},\overline{v}_{3})|=|\det(v_{1},\overline{v}_{3})|=h$, where $h$ is the height of $\overline{P}$ with respect to $e_{3}$, that is, the lattice distance between $e_{3}$ and $e_{4}$. It follows that the condition $w_{3}-w_{4}=|\det(v_{2},\overline{v}_{3})|$ is equivalent to $w_{3}-w_{4}=h$.
@@
Finally, we describe the geometric interpretation of pushing the 7-brane through the origin when $w_{3}-w_{4}=h$. Let $(\overline{Y},\overline{D},\overline{L})$ be the polarized toric surface corresponding to $\overline{W}$. The toric boundary divisor $\overline{D}$ has four irreducible toric divisors $\overline{D}_{1},\overline{D}_{2},\overline{D}_{3},\overline{D}_{4}$ corresponding to $e_{1},e_{2},e_{3},e_{4}$, and the log Calabi–Yau surface $Y$ is obtained from $\overline{Y}$ by blowing up a smooth point $x_{3}$ of the toric boundary on $\overline{D}_{3}$, see Figure 4.8. Moreover, denoting by $E$ the exceptional divisor, we have $L=\overline{L}\otimes\mathcal{O}(-w_{3}E)$. Denote by $E^{\prime}$ the strict transform of the fiber passing through $x_{3}$ of the $\mathbb{P}^{1}$-fibration on $\overline{Y}$ induced by the fact that both $\mathbb{R}_{\geq 0}v_{3}$ and $\mathbb{R}_{\neq 0}v_{3}$ are rays of the fan of $\overline{Y}$. One can check that $L\cdot H=h-w_{3}$. In particular, when $w_{3}-w_{4}=h$, we have $L\cdot H=-w_{4}$ and $L$ is not nef. When $W$ is part of a bigger consistent web, one can flop $E^{\prime}$, and the surface $Y$ becomes $Y^{\prime}=\mathbb{F}_{0}=\mathbb{P}^{1}\times\mathbb{P}^{1}$. In addition, the resulting line bundle $L^{\prime}$ satisfies $L^{\prime}\cdot D_{3}^{\prime}=L^{\prime}\cdot D_{4}^{\prime}=0$ and $L^{\prime}\cdot D_{1}^{\prime}=L^{\prime}\cdot D_{2}^{\prime}=h$, and so is nef. Contracting all curves with zero-intersection with $L^{\prime}$ produces a map $c:Y^{\prime}=\mathbb{P}^{1}\times\mathbb{P}^{1}\rightarrow\mathbb{P}^{1}$. The fact that the surface $Y^{\prime}$ is contracted onto $\mathbb{P}^{1}$ is the geometric realization of the fact that, after pushing the 7-brane across the origin, the 3-valent vertex at the origin of the web disappears and the remaining 5-branes form a straight line (taking into account the monodromy cut), which can be viewed as the fan of $\mathbb{P}^{1}$.
[…]
Figure 4.8. The log Calabi–Yau surface $(Y,D)$ and the contraction map $c:Y^{\prime}\rightarrow\mathbb{P}^{1}$.
@@
[…]
Figure 4.9. Examples of elementary lattice trapezoids.
@@
The three lattice trapezoids in Figure 4.9 are elementary.
@@
If $w_{3}-w_{4}=h$, then we have automatically $w_{1}=w_{2}$. This results from the fact that the web is balanced after pushing the 7-brane (see also [6, Proposition 2.12] for a direct proof). It follows that, in the Definition 4.20 of an elementary trapezoid, it is enough to require that one of the two remaining edges has lattice length one. It also follows that any lattice triangle with $w_{3}=h$, or any lattice trapezoid with $w_{3}-w_{4}=h$, can be naturally subdivided into a union of elementary triangles and elementary trapezoids, see Figure 4.10 for an example.
[…]
Figure 4.10. On the left, a non-elementary triangle. On the right, a subdivision into one elementary triangle and two elementary trapezoids.
@@ 4.6. Generic consistent webs and fully pushed 7-branes
There is a natural notion of perturbation for webs of 5-branes $\overline{W}$ (see [81, Definition 2.17]): we say that $\overline{W}^{\prime}$ is a perturbation of $\overline{W}$ if there exists a one-parameter family of webs of 5-branes $(\overline{W}^{\prime}_{t})_{1\geq t>0}$ with the same combinatorial type as $\overline{W}^{\prime}$, such that $\overline{W}^{\prime}_{1}=\overline{W}^{\prime}$, and $\overline{W}=\mathrm{lim}_{t\rightarrow 0}\overline{W}^{\prime}_{t}$, where the limit is taken in the sense of weighted graphs, that is, with weights adding up when edges coincide in the limit. We generalize this notion of perturbation for webs of 5-branes with 7-branes:
@@
A web of 5-branes $\overline{W}$ is generic if and only if the polyhedral decomposition to $\overline{W}$ is a triangulation into triangles of size one, that is, if $\overline{W}$ is 3-valent and every vertex of $\overline{W}$ is of multiplicity one (see [81, Proposition 2.19]).
@@
There exist generic webs of 5-branes with 7-branes with 3-valent vertices of multiplicity $>1$ or 4-valent vertices. For instance, 3-valent vertices dual to elementary triangles and 4-valent vertices dual to elementary trapezoids as in §4.5 are generic. Indeed, for a 3-valent vertex dual to an elementary triangle, using the notation of 4.5.1, the vertex adjacent to the $w_{3}$ 5-branes ending on the 7-brane in a non-trivial perturbation of the web is necessarily 3-valent with multiplicity $m^{\prime}<m=w_{3}$, and so would not be consistent.
@@
We first prove the result for $L^{2}>0$. An irreducible component of $((\mathcal{Y}^{\prime})^{\mathrm{pol}}_{0},(\mathcal{D}^{\prime})^{\mathrm{pol}}_{0},(\mathcal{L}^{\prime})^{\mathrm{pol}}_{0})$ is a polarized log Calabi–Yau surface $(Y_{i},D_{i},L_{i})$. If $(Y_{i},D_{i},L_{i})$ is not isomorphic to $(\mathbb{P}^{2},\partial\mathbb{P}^{2},\mathcal{O}(1))$, then it follows from [6, §2.4] that there exists a non-trivial degeneration of $(Y_{i},D_{i},L_{i})$ to a union of copies of $(\mathbb{P}^{2},\partial\mathbb{P}^{2},\mathcal{O}(1))$. The existence of such a degeneration implies that the web $W$ admits a non-trivial consistent perturbation, which contradicts the assumption that $W$ is generic. The case where $L^{2}=0$ can be treated similarly. ∎
@@
    For every $(p_{i},q_{i})$ 7-brane $x_{i}$, the half-line $x_{i}+\mathbb{R}_{<0}(p_{i},q_{i})$ does not intersect the web $W^{\prime}$, that is, the 7-brane $x_{i}$ cannot be pushed further without crossing $W^{\prime}$.

### v2

@@ 4.4. Flops and pushing 7-branes
obtained from $\pi:(\mathcal{Y},\mathcal{D},\mathcal{L})\rightarrow\mathbb{C}$ by a sequence of M1 flops. Indeed, pushing a 7-brane $x_{e,i}$ across $v$ in $W_{v}$ as in Case II-(4.15) with $w_{-e}\geq 1$ amounts to flopping the (1)(-1)-curve $E_{e,i}^{\prime}$ as in §4.1. On the other hand, pushing a 7-brane $x_{e,i}$ across $v$ in $W_{v}$ as in Case II-(4.15) with $w_{-e}=\mu=0$, or as in Case III-(4.16) does not change the degeneration, but only the toric model of the log Calabi–Yau surface $(Y_{0}^{v},\partial Y_{0}^{v},L_{0}^{v})$, as in the geometric description of Hanany–Witten moves in §4.1.
@@
Since $W$ is consistent, Lemma 4.13 implies the existence of a degeneration of log Calabi–Yau surfaces $\pi^{\prime}:(\mathcal{Y}^{\prime},\mathcal{D}^{\prime},\mathcal{L}^{\prime})\rightarrow\mathbb{C}$, obtained from $\pi:(\mathcal{Y},\mathcal{D},\mathcal{L})\rightarrow\mathbb{C}$ by a sequence of M1 flops, such that all line bundles $(L^{\prime})_{0}^{v}$ are nef. This sequence of M1 flops can be realized through successive pushing operations of 7-branes, yielding a modified web $W^{\prime}$. Indeed, as in the proof of Lemma 4.4, by [58, Proposition 3.27] (see also [24, Theorem 1]), any two toric models of a log Calabi–Yau surface are related by a sequence of corner blow-ups/blow-downs and elementary cluster transformations. Consequently, any interior (1)(-1)-curve can be obtained, after a finite sequence of pushing operations of 7-branes, as the (1)(-1)-curve $E_{e,i}^{\prime}$ corresponding to a pushed 7-brane $x_{e,i}^{\prime}$. ∎
@@ 4.5.1. Pushing a 7-brane through a 3-valent vertex
Let $W$ be the web of 5-branes with 7-branes obtained from $\overline{W}$ by requiring that all the $w_{3}$ 5-branes of direction $v_{3}$ end on the same 7-brane, see Figure 11.
Figure 11. Before and after pushing a 7-brane through a 3-valent vertex.
@@
It follows that we are in Cases I-II-III as in (4.14)-(4.15)-(4.16) if and only if $w_{3}^{2}>m$, $w_{3}^{2}=m$, and $w_{3}^{2}<m$ respectively. In particular, we are in Case II if and only if $w_{3}^{2}=m$, in which case the 7-brane can be pushed through the origin and is no longer attached to any 5-brane, as in Figure 11. The condition $w_{3}^{2}=m$ can be rephrased in terms of the lattice triangle $\overline{P}$ dual to $\overline{W}$. First note that the weights $w_{1},w_{2},w_{3}$ are the lattice lengths of the edges $e_{1},e_{2},e_{3}$ of $\overline{P}$ corresponding to the edges of $\overline{W}$. Denoting by $h$ the height of $\overline{P}$ with respect to $e_{3}$ the lattice distance between $e_{3}$ and the opposite vertex of $\overline{P}$, we have $m=w_{3}h$. It follows that the condition $w_{3}^{2}=m$ is equivalent to $w_{3}=h$.
We now describe the geometric interpretation of pushing the 7-brane through the origin $w_{3}=h$. Let $(\overline{Y},\overline{D},\overline{L})$ be the polarized toric surface corresponding to $\overline{W}$. The toric boundary divisor $\overline{D}$ has three irreducible toric divisors $\overline{D}_{1},\overline{D}_{2},\overline{D}_{3}$ corresponding to $e_{1},e_{2},e_{3}$, and the log Calabi–Yau surface $Y$ is obtained from $\overline{Y}$ by blowing up a smooth point $x_{3}$ of the toric boundary on $\overline{D}_{3}$, see Figure 12. Moreover, denoting by $E$ the exceptional divisor, we have $L=\overline{L}\otimes\mathcal{O}(-w_{3}E)$. Denote by $H$ the image in $Y$ of the strict transform $E^{\prime}$ of the fiber passing through $x_{3}$ of the $\mathbb{P}^{1}$-fibration on the toric blow-up of $\overline{Y}$ obtained by adding the ray $\mathbb{R}_{<0}v_{3}$ to the fan. One can check that $L\cdot H=h-w_{3}$. In particular, when $w_{3}=h$, we have $L\cdot H=0$ and so $L$ is nef. On the other hand, we have $H^{2}=0$ since $(E^{\prime})^{2}=-1$ and the toric divisor corresponding to $\mathbb{R}_{<0}v_{3}$ is contracted in $Y$. It follows that there is a fibration of $Y$ with fibers of class $H$. Finally, using that $L\cdot E=w_{3}>0$, one concludes that the map contracting all curves in $Y$ having zero-intersection with $L$ is a map $c:Y\rightarrow\mathbb{P}^{1}$. The fact that the surface $Y$ is contracted onto $\mathbb{P}^{1}$ is the geometric realization of the fact that, after pushing the 7-brane across the origin, the 3-valent vertex of the web disappears and the remaining 5-branes form a straight line (taking into account the monodromy cut), which can be viewed as the fan of $\mathbb{P}^{1}$.
Figure 12. The contraction map $c:Y\rightarrow\mathbb{P}^{1}$.
@@
Figure 13. Examples of lattice triangles with $w_{3}=h$.
@@
The left and middle lattice triangles in Figure 13 are elementary. The right triangle in Figure 13 satisfies $w_{3}=h$ but is not elementary.
@@ 4.5.2. Pushing a 7-brane through a 4-valent vertex
Let $\overline{W}$ be a web of 5-branes consisting of a single 4-valent vertex at the origin with edges of primitive directions of the form $\overline{v}_{1}$, $\overline{v}_{2}$, $\overline{v}_{3}$, $\overline{v}_{4}=-\overline{v}_{3}$. Let $w_{1},w_{2},w_{3},w_{4}$ be the weights of the edges, and denote $v_{i}=w_{i}\overline{v}_{i}$ for $1\leq i\leq 4$. By the balancing condition, we have $\sum_{i=1}^{4}v_{i}=0$. Let $W$ be the web of 5-branes with 7-branes obtained from $\overline{W}$ by requiring that all the 5-branes in direction $v_{3}$ all end on the same 7-brane, see Figure 14.
@@
Figure 14. Before and after pushing a 7-brane through a 4-valent vertex.
@@
Arguing as in §4.5.1, we obtain that, after moving the 7-brane through the origin, the number of 5-branes starting at the origin and ending on the 7-brane is $|\det(v_{2},\overline{v}_{3})|+w_{4}-w_{3}$. Note that, by the balancing condition, we have $|\det(v_{2},\overline{v}_{3})|=|\det(v_{1},\overline{v}_{3})|$. Hence, we are in Cases I-II-III as in (4.14)-(4.15)-(4.16) if and only if $w_{3}-w_{4}>|\det(v_{2},\overline{v}_{3})|$, $w_{3}-w_{4}=|\det(v_{2},\overline{v}_{3})|$, and $w_{3}-w_{4}<|\det(v_{2},\overline{v}_{3})|$ respectively. In particular, we are in Case II if and only if $w_{3}-w_{4}=|\det(v_{2},\overline{v}_{3})|$, in which case the 7-brane can be pushed through the origin and is no longer attached to 5-branes connected to the origin – see Figure 14. The condition $w_{3}-w_{4}=|\det(v_{2},\overline{v}_{3})|$ can be rephrased in terms of the lattice trapezoid $\overline{P}$ dual to $\overline{W}$. First note that the weights $w_{1},w_{2},w_{3},w_{4}$ are the weights of the edges $e_{1},e_{2},e_{3}$ of $\overline{P}$ corresponding to the edges of $\overline{W}$. Then, we have $|\det(v_{2},\overline{v}_{3})|=|\det(v_{1},\overline{v}_{3})|=h$, where $h$ is the height of $\overline{P}$ with respect to $e_{3}$, that is, the lattice distance between $e_{3}$ and $e_{4}$. It follows that the condition $w_{3}-w_{4}=|\det(v_{2},\overline{v}_{3})|$ is equivalent to $w_{3}-w_{4}=h$.
@@
Finally, we describe the geometric interpretation of pushing the 7-brane through the origin when $w_{3}-w_{4}=h$. Let $(\overline{Y},\overline{D},\overline{L})$ be the polarized toric surface corresponding to $\overline{W}$. The toric boundary divisor $\overline{D}$ has four irreducible toric divisors $\overline{D}_{1},\overline{D}_{2},\overline{D}_{3},\overline{D}_{4}$ corresponding to $e_{1},e_{2},e_{3},e_{4}$, and the log Calabi–Yau surface $Y$ is obtained from $\overline{Y}$ by blowing up a smooth point $x_{3}$ of the toric boundary on $\overline{D}_{3}$, see Figure 15. Moreover, denoting by $E$ the exceptional divisor, we have $L=\overline{L}\otimes\mathcal{O}(-w_{3}E)$. Denote by $E^{\prime}$ the strict transform of the fiber passing through $x_{3}$ of the $\mathbb{P}^{1}$-fibration on $\overline{Y}$ induced by the fact that both $\mathbb{R}_{\geq 0}v_{3}$ and $\mathbb{R}_{\leq 0}v_{3}$ are rays of the fan of $\overline{Y}$. One can check that $L\cdot H=h-w_{3}$. In particular, when $w_{3}-w_{4}=h$, we have $L\cdot H=-w_{4}$ and $L$ is not nef. When $W$ is part of a bigger consistent web, one can flop $E^{\prime}$, and the surface $Y$ becomes $Y^{\prime}=\mathbb{F}_{0}=\mathbb{P}^{1}\times\mathbb{P}^{1}$. In addition, the resulting line bundle $L^{\prime}$ satisfies $L^{\prime}\cdot D_{3}^{\prime}=L^{\prime}\cdot D_{4}^{\prime}=0$ and $L^{\prime}\cdot D_{1}^{\prime}=L^{\prime}\cdot D_{2}^{\prime}=h$, and so is nef. Contracting all curves with zero-intersection with $L^{\prime}$ produces a map $c:Y^{\prime}=\mathbb{P}^{1}\times\mathbb{P}^{1}\rightarrow\mathbb{P}^{1}$. The fact that the surface $Y^{\prime}$ is contracted onto $\mathbb{P}^{1}$ is the geometric realization of the fact that, after pushing the 7-brane across the origin, the 4-valent vertex at the origin of the web disappears and the remaining 5-branes form a straight line (taking into account the monodromy cut), which can be viewed as the fan of $\mathbb{P}^{1}$.
Figure 15. The log Calabi–Yau surface $(Y,D)$ and the contraction map $c:Y^{\prime}\rightarrow\mathbb{P}^{1}$.
@@
Figure 16. Examples of elementary lattice trapezoids.
@@
The three lattice trapezoids in Figure 16 are elementary.
@@
If $w_{3}-w_{4}=h$, then we have automatically $w_{1}=w_{2}$. This results from the fact that the web is balanced after pushing the 7-brane (see also [6, Proposition 2.12] for a direct proof). It follows that, in the Definition 4.20 of an elementary trapezoid, it is enough to require that one of the two remaining edges has lattice length one. It also follows that any lattice triangle with $w_{3}=h$, or any lattice trapezoid with $w_{3}-w_{4}=h$, can be naturally subdivided into a union of elementary triangles and elementary trapezoids, see Figure 17 for an example.
Figure 17. On the left, a non-elementary triangle. On the right, a subdivision into one elementary triangle and two elementary trapezoids.
@@ 4.6. Generic consistent webs and fully pushed 7-branes
There is a natural notion of perturbation for webs of 5-branes $\overline{W}$ (see [84, Definition 2.17]): we say that $\overline{W}^{\prime}$ is a perturbation of $\overline{W}$ if there exists a one-parameter family of webs of 5-branes $(\overline{W}^{\prime}_{t})_{1\geq t>0}$ with the same combinatorial type as $\overline{W}^{\prime}$, such that $\overline{W}^{\prime}_{1}=\overline{W}^{\prime}$, and $\overline{W}=\mathrm{lim}_{t\rightarrow 0}\overline{W}^{\prime}_{t}$, where the limit is taken in the sense of weighted graphs, that is, with weights adding up when edges coincide in the limit. We generalize this notion of perturbation for webs of 5-branes with 7-branes:
@@
A web of 5-branes $\overline{W}$ is generic if and only if the polyhedral decomposition to $\overline{W}$ is a triangulation into triangles of size one, that is, if $\overline{W}$ is 3-valent and every vertex of $\overline{W}$ is of multiplicity one (see [84, Proposition 2.19]).
@@
There exist generic webs of 5-branes with 7-branes with 3-valent vertices of multiplicity >1>1 or 4-valent vertices. For instance, 3-valent vertices dual to elementary triangles and 4-valent vertices dual to elementary trapezoids as in §4.5 are generic. Indeed, for a 3-valent vertex dual to an elementary triangle, using the notation of 4.5.1, the vertex adjacent to the $w_{3}$ 5-branes ending on the 7-brane in a non-trivial perturbation of the web is necessarily 3-valent with multiplicity $m^{\prime}<m=w_{3}$, and so would not be consistent.
@@
We first prove the result for $L^{2}>0$. An irreducible component of $((\mathcal{Y}^{\prime})^{\mathrm{pol}}_{0},(\mathcal{D}^{\prime})^{\mathrm{pol}}_{0},(\mathcal{L}^{\prime})^{\mathrm{pol}}_{0})$ is a polarized log Calabi–Yau surface $(Y_{i},D_{i},L_{i})$. If $(Y_{i},D_{i},L_{i})$ is not isomorphic to $(\mathbb{P}^{2},\partial\mathbb{P}^{2},\mathcal{O}(1))$, then, by [6, §2.4], there exists a non-trivial degeneration of $(Y_{i},D_{i},L_{i})$ to a union of copies of $(\mathbb{P}^{2},\partial\mathbb{P}^{2},\mathcal{O}(1))$. The existence of such a degeneration implies that the web $W$ admits a non-trivial consistent perturbation, contradicting the assumption that $W$ is generic. The case where $L^{2}=0$ can be treated similarly. ∎
@@
    For every $(p_{i},q_{i})$ 7-brane $x_{i}$, the half-line $x_{i}+\mathbb{R}_{<0}(p_{i},q_{i})$ intersects the web $W^{\prime}$, that is, the 7-brane $x_{i}$ cannot be pushed further without crossing $W^{\prime}$.